OpenStudy (anonymous):

1/2x - 5 + 1/2x - 10 = 15 (1/2 = fraction. just so you know.) please help. if i fail this class i'm fluttered.

OpenStudy (anonymous):

love how it white fluttered instead of the word i really wanted to use. you all know what i meant... f***ed

OpenStudy (callisto):

 which one do you mean in the question?

OpenStudy (anonymous):

the first one.

OpenStudy (callisto):

1/2x - 5 + 1/2x - 10 = 15 1/2x + 1/2x -5 - 10 = 15 x( 1/2 +1/2) -(5+10) = 15 x - 15 = 15 Can you get the answer?

OpenStudy (anonymous):

yes. thanks :D

OpenStudy (callisto):

and do you understand how to do it?

OpenStudy (anonymous):

lol yeaaaah.

OpenStudy (callisto):

That's good :)

OpenStudy (anonymous):

1/2x-5+1/2x-10=15 1/2x+1/2x-5=15 x=x/2=20 2x/2=20 2x=40 x/2=40/2 x=20